# Split Two-Periodic Aztec Diamond

### v3

@@ 4.1. Preliminaries of Saddle Functions
(70) $\Big{[}J(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\frac{1}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\varepsilon,2}(z)^{|N-(m^{\prime}+m)|}M_{\varepsilon}(z)$
Since we are assuming $x\neq\nicefrac{{1}}{{2}}$ it follows that $|N-(m+m^{\prime})|>0$, so we can deform the contour to a circle of radius one centered at the origin,
@@
(71) $\frac{1}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\varepsilon,2}(z)^{|N-(m^{\prime}+m)|}M_{\varepsilon}(z)$
Along this contour, $|r_{\varepsilon,2}(z)|<1$ and the matrix-valued function $M_{\varepsilon}$ is bounded and analytic, so we can bound the integral accordingly,
@@
(72) $\frac{1}{2\pi}\Bigg{|}\oint_{|z|=1}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\varepsilon,2}(z)^{N-(m+m^{\prime})}M_{\varepsilon}(z)\Bigg{|}\leq C|r_{\alpha,2}^{\text{max}}|^{N-(m+m^{\prime})}$
where $r_{\varepsilon,2}^{\text{max}}$ denotes the maximum value of $r_{\varepsilon,2}(z)$ on the unit circle and $C$ is some constant matrix. The result in Proposition 2.1 follows.
4.3. Asymptotics of $I^{\alpha}_{2,2}$
The goal of this section is to prove Proposition 2.2. As we can see from the statement, the proof depends on the relative magnitudes of $\alpha$ and $\beta$. This is because the branch cuts of the integral in $I_{2,2}^{\alpha}$ are different in the cases of $\beta<\alpha$ and $\alpha<\beta$. In the case of $\beta<\alpha$ the branch cuts are $(-\infty,-\alpha^{-2}]\cup[-\alpha^{2},0]$. Alternately, if $\alpha<\beta$ the branch cuts are $(-\infty,-\beta^{-2}]\cup[-\beta^{2},0]$. These branch cuts are depicted in Figure 12. We begin by proving the former case.
$-1/\alpha^{2}$$-\alpha^{2}$$-\beta^{2}$$-1/\beta^{2}$
$-1/\beta^{2}$$-\beta^{2}$$-\alpha^{2}$$-1/\alpha^{2}$
Figure 12. Branch cuts of the $w$-integral part of the interal in $I^{\alpha}_{2,2}$. The left hand image shows the branch cuts when $10<\beta<\alpha<$ 1, while the right hand image shows the branch cuts when $10<\alpha<\beta<$ 1. Note the left most branch cuts go off to infinity in both images.
4.3.1. Asymptotics on the $\alpha$-side when $\beta<\alpha$
Here we will prove the statement regarding $I^{\alpha}_{2,2}$ in Proposition 2.2 under the condition $10<\beta<\alpha<$ 1.
@@ 4.3.1. Asymptotics on the $\alpha$-side when $\beta<\alpha$
We start with the definition of $[I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)]_{i,j=0}^{1}$ presented in equation 28. Since the $w$-contours transverse the branch cuts, we can consider them as two separate contours, which we call $\gamma_{w,1}$ and $\gamma_{w,2}$. We also deform the $z$-contour to the steepest ascent contour through the saddle point $z_{2}^{*}$. We refer to this contour as $\gamma_{s,2}$ and it is described in Lemma 4.3. The resulting contours are depicted in Figure 13.
$\gamma_{w,1}$$\gamma_{w,2}$$\gamma_{s,2}$
Figure 13. Depiction of the contours $\gamma_{s,2}$, $\gamma_{w,1}$, and $\gamma_{w,2}$ in the case of $(x,y)\not\in\mathfrak{C}$.
We write,
(73) $\Big{[}I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{w,1}\cup\gamma_{w,2}}\oint_{\gamma_{s,2}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}\\ \times F_{\alpha,2}(z)\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,2}(z;x_{1},y_{1})\right]$
The function $\varphi_{\alpha,2}(z)$ is defined back in equation 27. We are now ready to utilize saddle point asymptotics. As stated in Lemma 4.2 the saddle function $\psi_{2}(z)$ always has one saddle point on the interval $(-\alpha^{-2},-\alpha^{2})$, which we refer to as $z_{2}^{*}$. The contours $\gamma_{w,1}$ and $\gamma_{w,2}$ not only lie away from the saddle point $z_{2}^{*}$, we can also guarantee that they both lie entirely inside of $\Omega^{-}_{2}$ (per Lemma 4.3). Thus
for all $w\in\gamma_{w,1}\cup\gamma_{w,2}$ and for all $z\in\gamma_{s,2}$. Moreover, Lemma 4.1 implies that as $w\to\infty$, $\operatorname{Re}\psi_{2}(w)\to-\infty$. Combining these to statement mean that the term
$\exp\Big{[}N\big{(}\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\big{)}\Big{]}$
is $O(e^{-cN})$ for some $c>0$, uniformly for $(z,w)\in\gamma_{z}\times\gamma_{w,1}\cup\gamma_{w,2}$. ∎
4.3.2. Asymptotics on the $\alpha$-side when $\beta>\alpha$
We will utilize the following lemma when proving the local asymptotics of $I^{\alpha}_{2,2}$ when $\beta>\alpha$. We state and prove the lemma below,
@@ 4.3.2. Asymptotics on the $\alpha$-side when $\beta>\alpha$
yielding the results stated in the lemma. ∎
The local asymptotic behavior on the $\alpha$-side of the interface depends on where the saddle point $z_{2}^{*}$ is located. Recall that $z_{2}^{*}$ may be anywhere along the interval $(-\alpha^{-2},-\alpha^{2})$ (we exclude the endpoints since these sit exactly on the interface). If $z^{*}_{2}\in(-\beta^{-2},-\beta^{2})$, then the proof is identical to the case in section 4.3.1 and the integral is $O(e^{-N})$. However, if the saddle point lies on one of the branch cuts, it is impossible for the $w$-contours be entirely inside of $\Omega^{-}_{2}$ as they were in Section 4.3.1. It is in these situations that our analysis differs.
@@
Before we detail the proof of Proposition 2.2 in this case, we should first explain when our analysis fall into this category. Suppose $z_{2}^{*}\in(-\alpha^{-2},-\beta^{-2}]$. Since $z_{2}^{*}$ is the only maximum of $\operatorname{Re}\psi_{\alpha,2}(z)$ (see Lemma 4.2) this implies that $\psi^{\prime}_{\alpha,2}(-\beta^{-2};x,y)<0$. This inequality gives the following relationship between $x$ and $y$,
@@
These inequalities describe the location of the strong-coupling region of the model. Examples of this region are given back in Figure 8. We are now prepared to consider the analysis of $[I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)]_{i,j=0}^{1}$ when $(x,y)$ satisfies one of the above inequalities.
$\gamma_{w,1}$$\gamma_{w,2}$$\gamma_{s,2}$
Figure 14. Depiction of the contours $\gamma_{s,2}$, $\gamma_{w,1}$, and $\gamma_{w,2}$ when $(x,y)\in\mathfrak{C}$.
Proof of Proposition 2.2 for $(x,y)\in\mathfrak{C}$.
Assume $z_{2}^{*}\in[-\beta^{2},-\alpha^{2})$. We start by deforming the contours in the same way we did for the case of $\beta<\alpha$. Figure 14 gives an example of the contour deformation in this case. We split up the integral by the $w$-contours,
(78) $\Big{[}I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{w,1}}\oint_{\gamma_{s,2}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}\\ \times F_{\alpha,2}(z)\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,2}(z;x_{1},y_{1})\right]\\ +\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{w,2}}\oint_{\gamma_{s,2}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}\\ \times F_{\alpha,2}(z)\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,2}(z;x_{1},y_{1})\right]$
Since $\gamma_{w,2}$ still lies entirely inside of $\Omega_{2}^{-}$, we know that
$\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{w,2}}\oint_{\gamma_{s,2}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,2}(z)\\ \times\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,2}(z;x_{1},y_{1})\right]$
@@
Now we separate the contour $\gamma_{w,1}$ into two pieces: $\gamma_{w,\uparrow}$, which is the half lying above the real axis, and $\gamma_{w,\downarrow}$, which is the half lying below the real axis. We deform both of these contours to the negative real axis, which is the contour of steepest descent through $z_{2}^{*}$. We will start by just considering the contour integral
(79) $\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{w,\uparrow}}\oint_{\gamma_{s,2}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,2}(z)\\ \times\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,2}(z;x_{1},y_{1})\right]$

### v4

@@ 4.2. Initial Deformation of the Contours of $I_{2,k}^{\alpha}$
(65) $\Big{[}I^{\alpha}_{2,k}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ =\frac{1}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\alpha,2}(z)^{\frac{N}{2}-m^{\prime}}r_{\alpha,k}(z)^{m-\frac{N}{2}}\frac{2}{1+2g_{\alpha,\beta}(z)}F_{\alpha,2}(z)F_{\beta,1}(z)F_{\alpha,1}(z)F_{\alpha,k}(z)\\ +\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,2}(w)^{\frac{N}{2}-m^{\prime}}r_{\alpha,k}(z)^{m-\frac{N}{2}}\\ \times\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)F_{\alpha,k}(z)\$
When $k=2$, the single contour integral above is exactly zero. This is due to the orthogonality of the $F_{\alpha}$-matrices. So we can write,
@@
(66) $\Big{[}I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ =\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\alpha,2}(w)^{\frac{N}{2}-m^{\prime}}r_{\alpha,2}(z)^{m-\frac{N}{2}}\\ \times\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)F_{\alpha,2}(z)\$
We may also write it in terms of the saddle functions,
@@
(67) $\Big{[}I^{\varepsilon}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ =\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{\text{Br}}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)F_{\varepsilon,2}(z)\\ \times\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,2}(z;x_{1},y_{1})\right]$
However when $k=1$, the single contour integral in equation (65) is not zero. We provide the following lemma which proves that it decays exponentially in $N$,
Lemma 4.5.
For $m,\,m^{\prime},\,\xi,$ and $\xi^{\prime}$ defined in equations (21) and (22) such that $x<\nicefrac{{1}}{{2}}$, $m,\,m^{\prime}<\nicefrac{{N}}{{2}}$ and $x_{1},\,x_{2}=o(N)$ ,
$\left|\frac{1}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\alpha,2}(z)^{N-(m^{\prime}+m)}\frac{2}{1+2g_{\alpha,\beta}(z)}F_{\alpha,2}(z)F_{\beta,1}(z)F_{\alpha,1}(z)\right|<Ce^{-cN}$
@@
Since there is no singularity at $z=1$, we deform the contour $\gamma_{0,1}$ to a circle of radius $1$ centered at the origin. We also write,
$M_{\alpha}(z)=\frac{2}{1+2g_{\alpha,\beta}(z)}F_{\alpha,2}(z)F_{\beta,1}(z)F_{\alpha,1}(z)$
for brevity. So the contour integral can be written as,
$\frac{1}{2\pi\textrm{i}}\oint_{|z|=1}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\alpha,2}(z)^{N-(m^{\prime}+m)}M_{\alpha}(z)$
On the unit circle, $M_{\alpha}(z)$ is analytic and so it is bounded. Additionally, $|r_{\alpha,2}(z)|<1$ on the unit circle. We may bound the integral accordingly,
(68) $\frac{1}{2\pi}\Bigg{|}\oint_{|z|=1}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\varepsilon,2}(z)^{N-(m+m^{\prime})}M_{\alpha}(z)\Bigg{|}\leq C|r_{\alpha,2}^{\text{max}}|^{N-(m+m^{\prime})}$
where $r_{\alpha,2}^{\text{max}}$ denotes the maximum value of $r_{\alpha,2}(z)$ on the unit circle and $C$ is some constant matrix. Since $x<\nicefrac{{1}}{{2}}$ and $x_{1},\,x_{2}=o(N)$, the exponent $N-(m+m^{\prime})=O(N)$ and is positive. The lemma follows. ∎
4.3. Local Asymptotics of $I^{\alpha}_{2,2}$
The goal of this section is to prove Proposition 2.1, which outlines the local asymptotics of the contour integral $I^{\alpha}_{2,2}$. The proposition is split into three cases and we will consider each of these cases separately. Before we prove the individual cases, we will first state and prove a lemma that will be used,
@@ 4.3. Local Asymptotics of $I^{\alpha}_{2,2}$
yielding the results stated. ∎
4.3.1. Case 1: The asymptotic coordinate is in the rough region
We start with the expression of $I^{\alpha}_{2,2}$ given in equation (67). From this form of the integral we will deform the contours to the contours of steepest descent/ascent detailed in Section 4.1 and Lemma 4.4. Recall that these contours are depicted in Figure 11. We express the contour integral in the following manner,
(70) $\Big{[}I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\oint_{\gamma_{R,D}}\oint_{\gamma_{R,A}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,2}(z)\\ \times\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,2}(z;x_{1},y_{1})\right]$
Now we can use saddle point methods to compute the leading order decay of this integral. Since there are two different saddle points, we must consider the contributions from the four combinations. [^12] We can let $w=z_{3}^{*}$ or $w=z_{4}^{*}$. The same applies for $z$, thus resulting in four combinations. The contributions where different saddle points are used for $w$ and $z$ are immediate. We get the leading order term,
(71) $\frac{F_{\alpha,2}(z_{l}^{*})F_{\beta,1}(z_{l}^{*})F_{\alpha,1}(z_{l}^{*})F_{\alpha,2}(z_{k}^{*})\exp\Big{[}N(\psi_{2}(z_{l}^{*})-\psi_{2}(z_{k}^{*}))+\varphi_{2}(z_{l}^{*})-\varphi_{2}(z_{k}^{*})+2\textrm{i}\,\theta_{lk}(x,y)\Big{]}}{\pi z_{k}^{*}(z_{k}^{*}-z_{l}^{*})(1+2g_{\alpha,\beta}(z_{l}^{*}))N\sqrt{|\psi_{2}^{\prime\prime}(z_{k}^{*})\psi_{2}^{\prime\prime}(z_{l}^{*})|}}$
where either $l=3$ and $k=4$ or $l=4$ and $k=3$. The function $\theta_{lk}(x,y)$ depends on the angle of the steepest ascent and descent contours around the saddle points. The term $\psi_{2}(z_{l}^{*})-\psi_{2}(z_{k}^{*})$ is purely imaginary, and thus does not contribute to the decay of the term. The absolute value of the above can be bounded by $CN^{-1}$, where $C$ is some constant matrix.
Now we want to consider the leading order contributions from the terms where we consider $z$ and $w$ at the same saddle point. We start by expanding the function $\psi_{2}(z)$ around $z=z_{k}^{*}$,
$\psi_{2}(z)=\psi_{2}(z_{k}^{*})+\psi_{2}^{\prime\prime}(z_{k}^{*})(z-z_{k}^{*})^{2}+O((z-z_{k}^{*})^{3})$
In addition, we make the following change of variables,
(72) $z-z_{k}^{*}=sN^{-1/2}\qquad\qquad w-z_{k}^{*}=rN^{-1/2}$
Combining this gives,
(73) $N\Big{[}\psi_{2}(w)-\psi_{2}(z)\Big{]}=\psi_{2}^{\prime\prime}(z_{k}^{*})(s^{2}-r^{2})+O(N^{-1/2})$
We also apply Lemma 4.6 to get,
(74) $\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)F_{\alpha,2}(z)}{1+2g_{\alpha,\beta}(w)}=\frac{2F_{\alpha,2}(z_{k}^{*})F_{\beta,1}(z_{k}^{*})F_{\alpha,1}(z_{k}^{*})F^{\prime}_{\alpha,2}(z_{k}^{*})}{1+2g_{\alpha,\beta}(z_{k}^{*})}N^{-1/2}(s-r)+O(N^{-1})$
Combining all this we get the leading order asymptotics,
(75) $\frac{1}{(2\pi)^{2}}\frac{2F_{\alpha,2}(z_{k}^{*})F_{\beta,1}(z_{k}^{*})F_{\alpha,1}(z_{k}^{*})F^{\prime}_{\alpha,2}(z_{k}^{*})}{z_{k}^{*}(1+2g_{\alpha,\beta}(z_{k}^{*}))}N^{-1}\int_{\Gamma_{r}}\int_{\Gamma_{s}}\exp\big{[}\psi_{2}^{\prime\prime}(z_{k}^{*})(s^{2}-r^{2})\big{]}\text{d}s\text{d}r$
Where $\Gamma_{r}$ and $\Gamma_{s}$ are lines through the origin whose angles depend on the angle of the contours of steepest descent and ascent, respectively. Moreover, as a result of choosing the steepest ascent/descent contours of $\psi_{2}$, we know the integral above converges. Combining the results from equations (71) and (75) we obtain the results presented in part (1) of Proposition 2.1.
4.3.2. Case 2: The asymptotic coordinate is in the smooth region, but not in the strong-coupling region.
We start with the definition of $I^{\alpha}_{2,2}$ presented in equation (67). We then deform $\gamma_{0,1}$ to the steepest ascent contour through the saddle point $z_{2}^{*}$. We refer to this contour as $\gamma_{s,2}$ and it is described in Lemma 4.3. The resulting contours are depicted in Figure 13.
$\gamma_{\text{Br},1}$$\gamma_{\text{Br},2}$$\gamma_{s,2}$
Figure 13. Depiction of the contours $\gamma_{s,2}$, $\gamma_{\text{Br},1}$, and $\gamma_{\text{Br},2}$ in the case of $(x,y)\in\mathfrak{S}$ and $(x,y)\not\in\mathfrak{C}$.
We write,
(76) $\Big{[}I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{\text{Br},1}\cup\gamma_{\text{Br},2}}\oint_{\gamma_{s,2}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}\\ \times F_{\alpha,2}(z)\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,2}(z;x_{1},y_{1})\right]$
By Lemma 4.2 and since we are considering $(x,y)\not\in\mathfrak{C}$, we know the saddle point, $z_{2}^{*}$, lies away from the branch cuts $(-\infty,-\beta^{-2})\cup(-\beta^{2},0]$ and somewhere on the interval $(-\beta^{-2},-\beta^{2})$. Corollary 4.1 guarantees the contours $\gamma_{\text{Br},1}$ and $\gamma_{\text{Br},2}$ can be deformed so that they lie entirely inside of $\Omega^{-}_{2}$. Thus
for all $w\in\gamma_{\text{Br},1}\cup\gamma_{\text{Br},2}$ and for all $z\in\gamma_{s,2}$. Moreover, Lemma 4.1 implies that as $w\to\infty$, $\operatorname{Re}\psi_{2}(w)\to-\infty$. Combining these to statement mean that the term
$\exp\Big{[}N\big{(}\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\big{)}\Big{]}$
is $O(e^{-cN})$ for some $c>0$, uniformly for $(z,w)\in\gamma_{z}\times\gamma_{\text{Br},1}\cup\gamma_{\text{Br},2}$.
4.3.3. Case 3: The asymptotic coordinate is in the smooth region and strong-coupling region.
The asymptotics become more complicated in the intersection of the smooth region and the strong-coupling region. This is because the saddle point $z_{2}^{*}$ now lies in the branch cuts (of the $w$ integral). While the deformation of the $z$-contour remains the same, it is no longer true that both the contours $\gamma_{\text{Br},1}$ and $\gamma_{\text{Br},2}$ will lie entirely inside of $\Omega_{2}^{-}$.
@@ 4.3.1. Case 1: The asymptotic coordinate is in the rough region
Before we detail the proof of case (3) in Proposition 2.1, we should first explain when our analysis fall into this category. Suppose $z_{2}^{*}\in(-\alpha^{-2},-\beta^{-2}]$, then $z_{4}^{*}>-1$. This implies that $\psi^{\prime}_{\alpha,2}(-\beta^{-2};x,y)<0$. This inequality gives the following relationship between $x$ and $y$,
@@
These inequalities describe the location of the strong-coupling region of the model. Examples of this region are given back in Figure 8. We are now prepared to consider the analysis of $I^{\alpha}_{2,2}$ for $(x,y)\in\mathfrak{S}\cap\mathfrak{C}$.
$\gamma_{\text{Br},1}$$\gamma_{\text{Br},2}$$\gamma_{s,2}$
Figure 14. Example of the contours $\gamma_{s,2}$, $\gamma_{\text{Br},1}$, and $\gamma_{\text{Br},2}$ when $(x,y)\in\mathfrak{S}\cap\mathfrak{C}$.
Proof of Proposition 2.1 (3).
We will assume $z_{2}^{*}\in[-\beta^{2},-\alpha^{2})$. The procedure is nearly identical in the case of $z_{2}^{*}\in(-\alpha^{-2},-\beta^{-2}]$, the only difference is that the $\gamma_{s,2}$ intersects $\gamma_{\text{Br},2}$ instead of $\gamma_{\text{Br},1}$.
We start by deforming the contours in the same way we did for the case (2) in Section 4.3.2. Figure 14 gives an example of the contour deformation in this case. We split up the integral according to the $w$-contours,
(80) $\Big{[}I^{\alpha}_{2,2}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}\\ =\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{\text{Br},1}}\oint_{\gamma_{s,2}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,2}(z)\\ \times\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,2}(z;x_{1},y_{1})\right]\\ +\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{\text{Br},2}}\oint_{\gamma_{s,2}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,2}(z)\\ \times\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,2}(z;x_{1},y_{1})\right]$
Since $\gamma_{\text{Br},2}$ can still be deformed to lie entirely inside of $\Omega_{2}^{-}$, we know that
$\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{\text{Br},2}}\oint_{\gamma_{s,2}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,2}(z)\\ \times\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,2}(z;x_{1},y_{1})\right]$
@@
Now we separate the contour $\gamma_{\text{Br},1}$ into two pieces: $\gamma_{\text{Br},\uparrow}$, which is the half lying above the real axis, and $\gamma_{\text{Br},\downarrow}$, which is the half lying below the real axis. We deform both of these contours to the negative real axis, which is the contour of steepest descent through $z_{2}^{*}$. We will start by just considering the contour integral
(81) $\frac{1}{(2\pi\textrm{i})^{2}}\oint_{\gamma_{\text{Br},\uparrow}}\oint_{\gamma_{s,2}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{2F_{\alpha,2}(w)F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}F_{\alpha,2}(z)\\ \times\exp\left[N\left(\psi_{2}(w;x,y)-\psi_{2}(z;x,y)\right)+\varphi_{\alpha,2}(w;x_{2},y_{2})-\varphi_{\alpha,2}(z;x_{1},y_{1})\right]$